OpenStudy (anonymous):

show that: sin(4θ=(cosθ)(4sinθ-8sin^3θ

OpenStudy (dumbcow):

\[\cos \theta(4\sin \theta -8\sin^{3} \theta)\] \[=4\sin \theta \cos \theta (1-2\sin^{2} \theta)\] cos double angle identity \[=4\sin \theta \cos \theta \cos (2\theta)\] sine double angle identity \[=2\sin (2\theta) \cos(2\theta)\] \[=\sin (4\theta)\]

OpenStudy (anonymous):

Thank you

OpenStudy (dumbcow):

yw :)